160 galons of 8% fat-milk is mixed with non-fat milk (0%) to make 2% fat-milk. How much non-fat milk is needed?
Answer by mananth(16949)

--------percent ---------------- quantity
fat milk 8 ---------------- 160 gallons
non fat milk 0 ---------------- x gallons
Mixture 2 ---------------- 160 + x gallons

8*160+0x=2(160+ x)
1280 +0x=320 +2x
0x+-2x = 320 -1280
-2x = -960
/ -2
x = 480
480 gallons of non fat milk